OpenStudy (anonymous):

Is 2 and 7/8s plus 5 and 1/2 = 8 and 3/8s?

OpenStudy (turingtest):

unmix your fractions: 23/8+11/2=67/8 get a common denominator (8): 23/8+44/8=67/8 67/8=67/8 so YES

OpenStudy (anonymous):

thank you

OpenStudy (turingtest):

yw, but medals suffice here :)